58% of all the town's residents own a dog and 62% own a cat. Of the dog owners 45% also own a cat. If a town resident is chosen at random find: (round to 4 decimal places where possible)
a. P(Own a Dog) =
b. P(Own a Cat) =
c. P(Own a Cat and a Dog) =
d. P(Own a Dog GIVEN Own a Cat) =

Answer by ikleyn(52879)

.
58% of all the town's residents own a dog and 62% own a cat. Of the dog owners 45% also own a cat.
If a town resident is chosen at random find: (round to 4 decimal places where possible)

a.  P(Own a Dog)                 = 58%,  or  0.58.

b.  P(Own a Cat)                 = 62%,  or  0.62.

c.  P(Own a Cat and a Dog)       = 0.58*0.45 = 0.261,  or  26.1%.

d.  P(Own a Dog GIVEN Own a Cat) = 0.261/0.62 = 0.4210,  or  42.1%.